# Adaptive Audio-Visual Speech Recognition via Matryoshka-Based Multimodal LLMs

### v1

@@ Abstract
Audio-Visual Speech Recognition (AVSR) leverages both audio and visual modalities to enhance speech recognition robustness, particularly in noisy environments. Recent advancements in Large Language Models (LLMs) have demonstrated their effectiveness in speech recognition, including AVSR. However, due to the significant length of speech representations, direct integration with LLMs imposes substantial computational costs. Prior approaches address this by compressing speech representations before feeding them into LLMs. However, higher compression ratios often lead to performance degradation, necessitating a trade-off between computational efficiency and recognition accuracy. To address this challenge, we propose Llama-MTSK, the first Matryoshka-based Multimodal LLM for AVSR, which enables flexible adaptation of the audio-visual token allocation based on specific computational constraints while preserving high performance. Our approach, inspired by Matryoshka Representation Learning, encodes audio-visual representations at multiple granularities within a single model, eliminating the need to train separate models for different compression levels. Moreover, to efficiently fine-tune the LLM, we introduce three LoRA-based Matryoshka strategies using global and scale-specific LoRA modules. Extensive evaluations on the two largest AVSR datasets demonstrate that Llama-MTSK achieves state-of-the-art results, matching or surpassing models trained independently at fixed compression levels.
@@
Adaptive Audio-Visual Speech Recognition
via Matryoshka-Based Multimodal LLMs
1 Introduction
@@
Audio-Visual Speech Recognition (AVSR) aims to improve the robustness of speech recognition systems by utilizing both audio and visual signals to recognize human speech. The correlation between audio and lip movements enables the model to focus on relevant speech content while discarding ambient or background noise. With the rising demand for robust speech recognition systems and the widespread availability of cameras (e.g., smartphones), numerous studies have explored advancements in AVSR technology. They have investigated different neural architectures (Dupont and Luettin, 2000; Noda et al., 2015; Afouras et al., 2018a; Petridis et al., 2018; Ma et al., 2021; Hong et al., 2022), training methods (Ma et al., 2023a; Hong et al., 2023), and methods using self-supervised pretraining (Shi et al., 2022; Haliassos et al., 2023, 2024b; Hsu and Shi, 2022; Haliassos et al., 2024a).
Collection and processing of the LRS2 and LRS3 datasets, and running of the Whisper (OpenAI) model was done by the Imperial College London authors on Imperial College London systems.
@@
Recently, with the growing popularity and versatility of Large Language Models (LLMs), new efforts have emerged to connect LLMs with speech modeling (Lakhotia et al., 2021; Huang et al., 2024; Park et al., 2024). Specifically, in Auditory Speech Recognition (ASR) and Visual Speech Recognition (VSR), researchers have demonstrated the possibility and effectiveness of LLMs in speech recognition (Chen et al., 2024; Hu et al., 2024b; Ma et al., 2024; Yu et al., 2024; Fathullah et al., 2024; Fang et al., 2024a; Lu et al., 2025; Tan et al., 2024; Yeo et al., 2024). By employing multi-modal speech information, a recent work proposes to adapt LLMs in AVSR as well (Llama-AVSR), attaining state-of-the-art recognition performances (Cappellazzo et al., 2025). A common focus of prior works is reducing the sequence length of speech representations before feeding them into the LLM. Since LLMs have a large number of parameters and speech sequences are much longer than text, directly using speech representations imposes a significant computational burden. At the same time, (Cappellazzo et al., 2025) demonstrate that there is a trade-off between how much we compress the audio-visual speech representations and performance: while higher compression ratios enhance computational efficiency, they lead to a degradation in performance. Therefore, a possible solution is training and distributing different models with compression ratios tailored to individual users’ computational resources.
@@
However, retraining existing models for different compression ratios, each requiring a distinct coarse-to-fine granularity, is time-consuming and impractical. For this reason, we propose to leverage the concept of Matryoshka Representation Learning (MRL) (Kusupati et al., 2022; Kudugunta et al., 2024; Nair et al., 2025) to encode audio-visual information at different granularities using a single model. This concept was recently explored in visual-linguistic understanding and reasoning tasks in (Cai et al., 2024; Hu et al., 2024a), demonstrating that Matryoshka-based large vision-language models can support multi-granular visual processing at inference while achieving performance parity with independently trained models for each compression rate.
@@
For our audio-visual setting, with the aspiration to flexibly decide between computational efficiency and performance at inference time within the same model, we propose Llama-Matryoshka (abbreviated as Llama-MTSK in the rest of the paper), a Matryoshka-based Multimodal LLM which caters to different demands based on specific requirements by training simultaneously audio-visual representations of different granularities. Llama-MTSK first produces audio and video tokens using pre-trained encoders, then reduces their length using average pooling or stacking compression methods at multiple compression rates. Then, unlike the previous works using MRL that directly fine-tune all the LLM’s parameters (Cai et al., 2024; Hu et al., 2024a), we propose three LoRA-based Matryoshka approaches (LoRA [Uncaptioned image] ) to parameter-efficiently fine-tune the LLM (i.e., Llama (Dubey et al., 2024)), which is responsible to generate the transcriptions given the audio-visual tokens and textual prompt. These approaches either employ a single global LoRA to learn audio-visual feature tokens at multiple scales (Multi-Scale LoRA [Uncaptioned image] ), or define multiple LoRAs, each of them focusing on scale-specific audio-visual information (Scale-Specific LoRA [Uncaptioned image] ), or a combination of both (Multi-Scale-Specific LoRA [Uncaptioned image] ). At inference, only the projector and LoRA modules associated with the desired compression rate are activated, ensuring both flexibility and efficiency. Our comprehensive experiments on the two largest AVSR datasets demonstrate that our three proposed methods achieve comparable or better performance than training separate models for each combination of audio-video compression rates. Overall, Llama-MTSK exhibits strong performance results, elastic inference, and computational efficiency under a single set of weights.
@@
2 Llama-MTSK
The objective of Llama-MTSK is to train an LLM (Llama-based in our setting) that captures audio and visual information at multiple scales, from coarse to fine, thus providing control over the audio-visual granularity during inference. Consequently, a single “universal” model allows us to dynamically adjust the performance-efficiency trade-off at inference time, according to specific needs (Cai et al., 2024; Hu et al., 2024a).
Llama-MTSK follows the structure of Llama-AVSR (Cappellazzo et al., 2025), the first Multimodal LLM (MLLM) tailored for audio-visual speech recognition, with ad-hoc modifications to support MRL (Kusupati et al., 2022). Llama-MTSK computes audio and video tokens via modality-specific pre-trained encoders, and then input them as prefix tokens to the LLM (together with the textual tokens). This approach, denoted as decoder-only, is adopted by several architectures due to its versatility and flexibility (Liu et al., 2023; Lin et al., 2024; Fang et al., 2024b; Fan et al., 2024; Zong et al., 2024; Zhang et al., 2025b; Lee et al., 2024; Fini et al., 2024; Li et al., 2024; Tong et al., 2024; Yao et al., 2024).
@@ 2 Llama-MTSK
2.1 Audio/Video Pre-Trained Encoders
We use pre-trained audio and video encoders to project the input audio and video data into two sets of audio and video tokens. We denote with $\mathbf{X}^{\mathsf{A}}\in\mathbb{R}^{N_{\mathsf{A}}\times d_{\mathsf{A}}}$ and $\mathbf{X}^{\mathsf{V}}\in\mathbb{R}^{N_{\mathsf{V}}\times d_{\mathsf{V}}}$ the audio and video token sequences, respectively, where $N_{\mathsf{A}}$/$N_{\mathsf{V}}$ is the number of audio/video tokens, and $d_{\mathsf{A}}$/$d_{\mathsf{V}}$ is the audio/video token dimension. The pre-trained encoders are maintained frozen during the training stage ( [Uncaptioned image] in Figure 1).
2.2 Audio-Visual Compression and Projection
Since the dimensions of audio and video tokens often differ from that of the textual tokens, MLLMs include a projection layer that maps audio and video tokens into the LLM embedding space. It is common to employ either linear projectors (Liu et al., 2023; Luo et al., 2024; Yao et al., 2024; Li et al., 2024; Liu et al., 2024b; Zhang et al., 2025a) or abstractors (e.g., Q-Former, resampler) (Zhu et al., 2023; Li et al., 2023; Cha et al., 2024). In our setting, following (Cappellazzo et al., 2025), we use a two-layer MLP projector.
In addition to this, since the LLM predominantly accounts for the entire computation and memory consumption of the MLLM, it is customary to compress the number of multimodal tokens (in our case audio-visual tokens) by a specific factor in order to find the optimal balance in terms of efficiency and accuracy. For example, (Cappellazzo et al., 2025; Fang et al., 2024a; Ma et al., 2024; Fathullah et al., 2024) stack multiple consecutive tokens along the token hidden dimension to reduce the number of tokens, whereas other methods rely on the Q-Former architecture (Li et al., 2023) using a fixed number of query tokens (Tang et al., 2023; Yu et al., 2024; Zhang et al., 2025b; Cha et al., 2024). However, all these methods need to decide the compression rate to apply beforehand, which means they generate outputs of a single, predetermined length, lacking the ability to modulate the final sequence length. This constraint limits the ability to balance information density and computational efficiency, particularly in resource-constrained deployment scenarios. Alternatively, one could train a separate model for each desired compression rate, but this approach can be time-consuming and cumbersome in practice.
@@ 2.2 Audio-Visual Compression and Projection
In contrast, we propose to compress the audio and video tokens using multiple compression rates, leading to token sequences at multiple scales, and thus different granularities. We explore two different compression methods to reduce the token sequence length: 1) average pooling, and 2) hidden size stacking, where multiple consecutive frames are stacked along the token hidden dimension. Therefore, we decide beforehand a range of G audio compression rates {$a_{1},a_{2},\cdots,a_{\texttt{G}}$} and $T$ video compression rates {$v_{1},v_{2},\cdots,v_{\texttt{T}}$}. We gradually increase the compression rates (i.e., $a_{i+1}>a_{i}$, $i=1,\cdots,\texttt{G}$). With $a_{i}$ we refer both to the compression rate and the corresponding scale interchangeably (e.g., if $a_{i}$ = $4$, then the corresponding sequence would have $\lfloor\frac{N_{\mathsf{A}}}{4}\rfloor$ tokens). We then compress the audio and video tokens using the chosen rates, producing token sequences at multiple scales: [$\mathbf{X}^{\mathsf{A}}_{a_{1}},\mathbf{X}^{\mathsf{A}}_{a_{2}},\cdots,\mathbf{X}^{\mathsf{A}}_{a_{\texttt{G}}}$] and [$\mathbf{X}^{\mathsf{V}}_{v_{1}},\mathbf{X}^{\mathsf{V}}_{v_{2}},\cdots,\mathbf{X}^{\mathsf{V}}_{v_{\texttt{T}}}$].
@@
2.3 LLM Adaptation via LoRA [Uncaptioned image]
The LLM is responsible for generating the corresponding ASR transcription in an auto-regressive fashion given the audio, video, and textual tokens. We define $\mathbf{X}^{\mathsf{AV}}_{ij}$ as the concatenation of audio and video tokens with audio and video compression rates of $a_{i}$ and $v_{j}$, and the prompt textual tokens $\mathbf{X}^{P}$: $\mathbf{X}^{\mathsf{AV}}_{ij}=[\mathbf{X}^{\mathsf{A}}_{a_{i}},\mathbf{X}^{\mathsf{V}}_{v_{j}},\mathbf{X}^{P}]$. To parameter-efficiently align the LLM with the multimodal inputs, we use LoRA modules (Hu et al., 2021) to adapt the query and value projection matrices of each layer. In our setting, the LLM is trained on multiple audio-visual tokens with different scales. We investigate three different strategies to efficiently fine-tune LLM’s pre-trained matrices via LoRA approximation under a MRL setting: 1) Multi-Scale LoRA Matryoshka (MS LoRA [Uncaptioned image] ), 2) Specific-Scale LoRA Matryoshka (SS LoRA [Uncaptioned image] ), and 3) Multi-Specific-Scale LoRA Matryoshka (MSS LoRA [Uncaptioned image] ). These three methods are illustrated in detail in Figure 2.
@@ 2.3 LLM Adaptation via LoRA [Uncaptioned image]
[…]
@@
where $W_{\texttt{SS}}^{ij}$ is the LoRA decomposition matrix defined for the $i$-th audio scale and $j$-th video scale, and it is defined as $W_{MS}$. As we explain in subsection 2.4, while all the LoRA modules are used during the training stage, at inference we only activate one LoRA module, corresponding to the selected audio and video scales.
@@
Regardless of the LoRA [Uncaptioned image] fine-tuning approach we employ, Llama-MTSK is trained by averaging the auto-regressive next token prediction loss for each audio-visual scale $ij$ for each input data. The LLM predicts the response $\mathbf{Y}=\{y_{l}\}_{l=1}^{L}$ conditioned on the multimodal input tokens, where $L$ represents the number of tokens of the ground truth transcription to be generated. Accordingly, for each Matryoshka audio-visual representation $\mathbf{X}^{\mathsf{AV}}_{ij}$, the probability of the target $\mathbf{Y}$ is computed by:
@@
where $y_{<l}$ is the generated output sequence up to token $l-1$, and $\theta$ is the trainable parameters, which comprises the projection layers and the LoRA [Uncaptioned image] modules according to the LoRA [Uncaptioned image] fine-tuning approach used.
@@
2.4 Llama-MTSK: Training vs Inference
During training, Llama-MTSK learns multiple sets of audio-visual tokens, each progressively incorporating more details as the scale increases. To do so, the LLM processes all the multi-scale audio-visual tokens and concurrently optimize over them using Eq. 5. This means that all the projectors and LoRA [Uncaptioned image] modules are involved. Instead, at inference time, for each input data, we choose a specific audio-visual scale and we activate only the projector and LoRA module associated with it. This is equivalent to one single Llama-AVSR model trained on the specific scale. This principle is similar to the behaviour of Mixture of Experts-based models (Shazeer et al., 2017; Fedus et al., 2022; Zoph et al., 2022; Mustafa et al., 2022; Puigcerver et al., 2023; Cappellazzo et al., 2024a; Jiang et al., 2024; Muennighoff et al., 2024), which at inference time only activate a small subset of the available experts (in our case the “experts” are the projectors and LoRA [Uncaptioned image] modules). Figure 1 depicts a schematic comparison of Llama-MTSK training and inference processes.
3 Experiments and Results
3.1 Implementation Details
Datasets. We train and evaluate Llama-MTSK on LRS2 (Son Chung et al., 2017) and LRS3 (Afouras et al., 2018b), the two largest publicly available datasets for audio-visual speech recognition. LRS2 includes $225$ hours of video clips from BBC programs. LRS3 contains $433$ hours of transcribed English video clips from TED talks.
Pre-Processing. We follow (Ma et al., 2023b; Cappellazzo et al., 2025) for the pre-processing of the datasets. For the video modality, we crop the mouth region of interests (ROIs) through a bounding box of $96$ $96$. Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set. Audio data only undergo z-normalisation per utterance.

### v2

@@ Abstract
Audio-Visual Speech Recognition (AVSR) leverages audio and visual modalities to improve robustness in noisy environments. Recent advances in Large Language Models (LLMs) show strong performance in speech recognition, including AVSR. However, the long speech representations lead to high computational costs for LLMs. Prior methods compress inputs before feeding them to LLMs, but high compression often harms accuracy. To address this, we propose Llama-MTSK, the first Matryoshka-based Multimodal LLM for AVSR, which flexibly adapts audio-visual token allocation under varying compute constraints. Inspired by Matryoshka Representation Learning, our model encodes representations at multiple granularities with a single architecture, avoiding the need for separate models. For efficient fine-tuning, we introduce three LoRA-based strategies using global and scale-specific modules. Evaluations on major AVSR datasets show Llama-MTSK matches or outperforms models trained at fixed compression levels.
Index Terms:
@@
Audio-Visual Speech Recognition, Matryoshka Representation Learning, Elastic Inference
I Introduction
@@
Audio-Visual Speech Recognition (AVSR) aims to improve the robustness of speech recognition systems by utilizing both audio and visual signals to recognize human speech. The correlation between audio and lip movements enables the model to focus on relevant speech content while discarding ambient or background noise. With the rising demand for robust speech recognition systems and the widespread availability of cameras (e.g., smartphones), numerous studies have explored advancements in AVSR technology. They have investigated different neural architectures [1, 2, 3, 4, 5, 6], training methods [7, 8], and methods using self-supervised pretraining [9, 10, 11, 12, 13].
@@
Recently, with the growing popularity and versatility of Large Language Models (LLMs), new efforts have emerged to connect LLMs with speech modeling [14, 15, 16]. Specifically, in Auditory Speech Recognition (ASR) and Visual Speech Recognition (VSR), researchers have demonstrated the possibility and effectiveness of LLMs in speech recognition [17, 18, 19, 20, 21, 22, 23, 24, 25]. By employing multi-modal speech information, recent works propose to adapt LLMs in AVSR as well, attaining state-of-the-art recognition performances [26, 27]. A common focus of prior works is reducing the sequence length of speech representations before feeding them into the LLM. Since LLMs have a large number of parameters and speech sequences are much longer than text, directly using speech representations imposes a significant computational burden. At the same time, [26] demonstrate that there is a trade-off between how much we compress the audio-visual speech representations and performance: while higher compression ratios enhance computational efficiency, they lead to a degradation in performance. Therefore, a possible solution is training and distributing different models with compression ratios tailored to individual users’ computational resources.
@@
However, retraining existing models for different compression ratios, each requiring a distinct coarse-to-fine granularity, is time-consuming and impractical. For this reason, we propose to leverage the concept of Matryoshka Representation Learning (MRL) [28, 29, 30] to encode audio-visual information at different granularities using a single model. This concept was recently explored in visual-linguistic understanding and reasoning tasks in [31, 32], demonstrating that Matryoshka-based large vision-language models can support multi-granular visual processing at inference while achieving performance parity with independently trained models for each compression rate.
@@
For our audio-visual setting, with the aspiration to flexibly decide between computational efficiency and performance at inference time within the same model, we propose Llama-Matryoshka (abbreviated as Llama-MTSK in the rest of the paper), a Matryoshka-based Multimodal LLM which caters to different demands based on specific requirements by training simultaneously audio-visual representations of different granularities. Llama-MTSK first produces audio and video tokens using pre-trained encoders, then reduces their length using average pooling or stacking compression methods at multiple compression rates. Then, unlike the previous works using MRL that directly fine-tune all the LLM’s parameters [31, 32], we propose three LoRA-based Matryoshka approaches (LoRA [Uncaptioned image] ) to parameter-efficiently fine-tune the LLM (i.e., Llama [33]), which is responsible to generate the transcriptions given the audio-visual tokens and textual prompt. These approaches either employ a single global LoRA to learn audio-visual feature tokens at multiple scales (Multi-Scale LoRA [Uncaptioned image] ), or define multiple LoRAs, each of them focusing on scale-specific audio-visual information (Scale-Specific LoRA [Uncaptioned image] ), or a combination of both (Multi-Scale-Specific LoRA [Uncaptioned image] ). At inference, only the projector and LoRA modules associated with the desired compression rate are activated, ensuring both flexibility and efficiency. Our comprehensive experiments on the two largest AVSR datasets demonstrate that our three proposed methods achieve comparable or better performance than training separate models for each combination of audio-video compression rates. Overall, Llama-MTSK exhibits strong performance results, elastic inference, and computational efficiency under a single set of weights.
@@
II Llama-MTSK
The objective of Llama-MTSK is to train an LLM (Llama-based in our setting) that captures audio and visual information at multiple scales, from coarse to fine, thus providing control over the audio-visual granularity during inference. Consequently, a single “universal” model allows us to dynamically adjust the performance-efficiency trade-off at inference time, according to specific needs [31, 32].
Llama-MTSK follows the structure of Llama-AVSR [26], the first Multimodal LLM (MLLM) tailored for audio-visual speech recognition, with ad-hoc modifications to support MRL [28]. Llama-MTSK computes audio and video tokens via modality-specific pre-trained encoders, and then input them as prefix tokens to the LLM (together with the textual tokens). This approach, denoted as decoder-only, is adopted by several architectures due to its versatility and flexibility [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@ II Llama-MTSK
II-A Audio/Video Pre-Trained Encoders
We use pre-trained audio and video encoders to project the input audio and video data into two sets of audio and video tokens. We denote with $\mathbf{X}^{\mathsf{A}}\in\mathbb{R}^{N_{\mathsf{A}}\times d_{\mathsf{A}}}$ and $\mathbf{X}^{\mathsf{V}}\in\mathbb{R}^{N_{\mathsf{V}}\times d_{\mathsf{V}}}$ the audio and video token sequences, respectively, where $N_{\mathsf{A}}$$N_{\mathsf{V}}$ is the number of audio/video tokens, and $d_{\mathsf{A}}$$d_{\mathsf{V}}$ is the audio/video token dimension. The pre-trained encoders are maintained frozen during the training stage ( [Uncaptioned image] in Figure 1).
II-B Audio-Visual Compression and Projection
Since the dimensions of audio and video tokens often differ from that of the textual tokens, MLLMs include a projection layer that maps audio and video tokens into the LLM embedding space. It is common to employ either linear projectors [34, 45, 44, 42, 46, 47] or abstractors (e.g., Q-Former, resampler) [48, 49, 50]. In our setting, following [26], we use a two-layer MLP projector.
In addition to this, since the LLM predominantly accounts for the entire computation and memory consumption of the MLLM, it is customary to compress the number of multimodal tokens (in our case audio-visual tokens) by a specific factor in order to find the optimal balance in terms of efficiency and accuracy. For example, [26, 22, 19, 21] stack multiple consecutive tokens along the token hidden dimension to reduce the number of tokens, whereas other methods rely on the Q-Former architecture [49] using a fixed number of query tokens [51, 20, 39, 50]. However, all these methods need to decide the compression rate to apply beforehand, which means they generate outputs of a single, predetermined length, lacking the ability to modulate the final sequence length. This constraint limits the ability to balance information density and computational efficiency, particularly in resource-constrained deployment scenarios. Alternatively, one could train a separate model for each desired compression rate, but this approach can be time-consuming and cumbersome in practice.
@@ II-B Audio-Visual Compression and Projection
In contrast, we propose to compress the audio and video tokens using multiple compression rates, leading to token sequences at multiple scales, and thus different granularities. We explore two different compression methods to reduce the token sequence length: 1) average pooling, and 2) hidden size stacking, where multiple consecutive frames are stacked along the token hidden dimension. Therefore, we decide beforehand a range of G audio compression rates $a_{1},a_{2},\cdots,a_{\texttt{G}}$} and $T$ video compression rates $v_{1},v_{2},\cdots,v_{\texttt{T}}$}. We gradually increase the compression rates (i.e., $a_{i+1}>a_{i}$, $i=1,\cdots,\texttt{G}$). With $a_{i}$ we refer both to the compression rate and the corresponding scale interchangeably (e.g., if $a_{i}$ = $4$, then the corresponding sequence would have $\lfloor\frac{N_{\mathsf{A}}}{4}\rfloor$ tokens). We then compress the audio and video tokens using the chosen rates, producing token sequences at multiple scales: $\mathbf{X}^{\mathsf{A}}_{a_{1}},\mathbf{X}^{\mathsf{A}}_{a_{2}},\cdots,\mathbf{X}^{\mathsf{A}}_{a_{\texttt{G}}}$] and $\mathbf{X}^{\mathsf{V}}_{v_{1}},\mathbf{X}^{\mathsf{V}}_{v_{2}},\cdots,\mathbf{X}^{\mathsf{V}}_{v_{\texttt{T}}}$].
@@
II-C LLM Adaptation via LoRA [Uncaptioned image]
The LLM is responsible for generating the corresponding ASR transcription in an auto-regressive fashion given the audio, video, and textual tokens. We define $\mathbf{X}^{\mathsf{AV}}_{ij}$ as the concatenation of audio and video tokens with audio and video compression rates of $a_{i}$ and $v_{j}$, and the prompt textual tokens $\mathbf{X}^{P}$: $\mathbf{X}^{\mathsf{AV}}_{ij}=[\mathbf{X}^{\mathsf{A}}_{a_{i}},\mathbf{X}^{\mathsf{V}}_{v_{j}},\mathbf{X}^{P}]$. To parameter-efficiently align the LLM with the multimodal inputs, we use LoRA modules [52] to adapt the query and value projection matrices of each layer. In our setting, the LLM is trained on multiple audio-visual tokens with different scales. We investigate three different strategies to efficiently fine-tune LLM’s pre-trained matrices via LoRA approximation under a MRL setting: 1) Multi-Scale LoRA Matryoshka (MS LoRA [Uncaptioned image] ), 2) Specific-Scale LoRA Matryoshka (SS LoRA [Uncaptioned image] ), and 3) Multi-Specific-Scale LoRA Matryoshka (MSS LoRA [Uncaptioned image] ). These three methods are illustrated in detail in Figure 2.
@@ II-C LLM Adaptation via LoRA [Uncaptioned image]
In contraposition to MS LoRA [Uncaptioned image] , we propose to learn “expert” LoRA modules, which specialize to each scale. We call this approach Specific-Scale (SS) LoRA [Uncaptioned image] . Therefore, we define $\texttt{G}\cdot\texttt{T}$ LoRA modules, one for each audio-visual scale. We compute the projection output as follows:
@@
where $W_{\texttt{SS}}^{ij}$ is the LoRA decomposition matrix defined for the $i$-th audio scale and $j$-th video scale, and it is defined as $W_{MS}$. As we explain in subsection II-D, while all the LoRA modules are used during the training stage, at inference we only activate one LoRA module, corresponding to the selected audio and video scales.
@@
Regardless of the LoRA [Uncaptioned image] fine-tuning approach we employ, Llama-MTSK is trained by averaging the auto-regressive next token prediction loss for each audio-visual scale $ij$ for each input data. The LLM predicts the response $\mathbf{Y}=\{y_{l}\}_{l=1}^{L}$ conditioned on the multimodal input tokens, where $L$ is the number of tokens of the ground truth transcription to generate. Accordingly, for each Matryoshka audio-visual representation $\mathbf{X}^{\mathsf{AV}}_{ij}$, the probability of the target $\mathbf{Y}$ is computed by:
@@
where $y_{<l}$ is the generated output sequence up to token l1l-1, and $\theta$ is the trainable parameters, which comprises the projection layers and the LoRA [Uncaptioned image] modules according to the LoRA [Uncaptioned image] fine-tuning approach used.
@@
II-D Llama-MTSK: Training vs Inference
During training, Llama-MTSK learns multiple sets of audio-visual tokens, each progressively incorporating more details as the scale increases. To do so, the LLM processes all the multi-scale audio-visual tokens and concurrently optimize over them using Eq. 5. This means that all the projectors and LoRA [Uncaptioned image] modules are involved. Instead, at inference time, for each input data, we choose a specific audio-visual scale and we activate only the projector and LoRA module associated with it. This is equivalent to one single Llama-AVSR model trained on the specific scale. This principle is similar to the behaviour of Mixture of Experts-based models [53, 54, 55, 56, 57, 58, 59, 60], which at inference time only activate a small subset of the available experts (in our case the “experts” are the projectors and LoRA [Uncaptioned image] modules). Figure 1 depicts a schematic comparison of Llama-MTSK training and inference processes.
III Experiments and Results
III-A Implementation Details
Datasets. We train and evaluate Llama-MTSK on LRS2 [61] and LRS3 [62], the two largest publicly available datasets for audio-visual speech recognition. LRS2 includes $225$ hours of video clips from BBC programs. LRS3 contains $433$ hours of transcribed English video clips from TED talks.
Pre-Processing. We follow [63, 26] for the pre-processing of the datasets. For the video modality, we crop the mouth region of interests (ROIs) through a bounding box of $96$ × $96$. Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set. Audio data only undergo z-normalisation per utterance.